# Truth Knows No Language: Evaluating Truthfulness Beyond English

### v1

@@ 5 Discussion
The results of the manually translated extension to TruthfulQA revealed a correlation between language-resource availability and model truthfulness. Thus, LLMs demonstrated optimal truthfulness metrics in English, which has the highest volume of training data, while performing substantially lower in Basque, the language with the most limited resources. Detailed examination of response patterns from Gemma 2 27B, the model achieving the best overall results, indicated that English-language outputs consistently exhibited better content moderation, longer response length, and more comprehensive explanatory content. However, this level of sophistication in the responses was not replicated in other languages. Furthermore, several base models displayed fundamental comprehension deficiencies when processing Basque-language questions, suggesting significant limitations in low-resource language processing capabilities.
@@
Comparison with manual annotation demonstrated that our Judge-LLM obtained higher correlation scores with respect to human judgments. In fact, the IAA between the Judge-LLM and the manual evaluation is quite high, which demonstrates the superiority of using LLM-as-a-Judge over multiple-choice (MC2) to evaluate truthfulness.
In order to investigate potential self-evaluation bias in the LLM-as-a-Judge evaluation, we conducted a comparative analysis between judges from different model families. Specifically, we trained an additional Judge-LLM using Llama 3.1 8B, [^7] Correlations of this model are also reported in Table 3. maintaining identical training parameters and protocols as our primary Judge-LLM (Gemma 2 9b instruct). The obtained evaluation scores (see Table 8 in Appendix D) from both judges showed no significant differences in their assessment patterns, whether evaluating responses from their own model family or from others. Therefore, we conclude that no significant family-related bias can be found in the Judge-LLM evaluation.
@@
In contrast with Lin et al. (2022) and Aula-Blasco et al. (2025), we found that larger models in the same model family tend to outperform their smaller counterparts. This could be partially explained by the post-training alignment and larger size of the models we experiment with. These results are consistent with those obtained for other modern evaluations such as SimpleQA Wei et al. (2024). This pattern is observed for both base and instruct models.
Time and contextual-dependency crucial to evaluate truthfulness.
Figure 2 has shown that all models answer remarkably more truthfully to questions about universal topics in all languages, with scores around 90% accuracy. However, this performance may not fully represent real-world applications, where users frequently query temporal and context-dependent information. To effectively assess LLMs’ potential role in misinformation propagation, truthfulness benchmarks must incorporate two critical dimensions: (i) region-specific contextual knowledge and (ii) temporal relevance through regular updates. Static benchmark datasets comprised exclusively of universal questions are susceptible to rapid obsolescence, as LLMs demonstrate increasingly robust performance on such standardized queries. Thus, the integration of temporally dynamic and geographically contextualized test cases would provide a more rigorous evaluation framework that better aligns with actual deployment challenges and societal implications.
@@ 6 Related Work
VeritasQA is a multilingual dataset to evaluate truthfulness in LLMs, currently available in English, Spanish, Catalan, and Galician. The dataset is designed to be transferable across languages, context-independent, and temporally stable. The empirical evidence presented in Figure 2 indicates that LLMs are approaching performance saturation on datasets such as VeritasQA that exclusively test universal knowledge.
@@ 7 Conclusion
The challenge of evaluating generative tasks automatically is another significant limitation. On the one hand, a manual evaluation was deemed impractical due to the significant burden it would impose on resources and scalability. On the other hand, although automated methods like LLM-as-judge work well, they still fall short of capturing the accuracy and subtle nuances that manual evaluations can achieve — especially as models continue to improve.
Ours is a static dataset. And, as highlighted in this work, an important limitation of static truthfulness benchmarks is their lack of local knowledge and dynamism. For these benchmarks to effectively prevent the spread of falsehoods, they must be informed by local contexts and be continuously updated. By incorporating locally relevant and timely questions, benchmarks can more effectively reflect and address real-world challenges.
@@ Acknowledgments
We are also thankful to the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and by FEDER, EU; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iii) NÓS-ILENIA (2022/TL22/0021533). Also, Xunta de Galicia: Centro de investigación de Galicia accreditation 2024-2027 ED431G-2023/04. Blanca Calvo Figueras is supported by the UPV/EHU PIF22/84 predoc grant. Julen Etxaniz holds a PhD grant from the Basque Government (PRE_2024_2_0028).
@@ Appendix C Translation Metrics
Table 7: Evaluation of the machine translated TruthfulQA dataset using the human translations as reference.
Appendix D Comparing the results of two Judge-LLMs from different families
@@
Appendix E Comparison between averaged results of Instruct models and Base models
Figure 5: Average performance of instruct and base models per language. The languages are ordered from higher to lower resourced.

### v2

@@ 5 Discussion
The results of the manually translated extension to TruthfulQA revealed a correlation between textual resource availability and model truthfulness. Thus, LLMs demonstrated optimal truthfulness metrics in English, which has the highest volume of training data, while performing substantially lower in Basque, the language with the most limited resources. Detailed examination of response patterns from Gemma 2 27B, the model achieving the best overall results, indicated that English-language outputs consistently exhibited better content moderation, longer response length, and more comprehensive explanatory content (see an example in Annex F.) However, this level of sophistication in the responses was not replicated in other languages. Furthermore, several base models displayed comprehension deficiencies when processing Basque-language questions, suggesting significant limitations in low-resource language processing capabilities.
@@
Comparison with manual annotation demonstrated that our Judge-LLM obtained a higher correlation with respect to human judgments. In fact, the IAA between the Judge-LLM and the manual evaluation is quite high, which demonstrates the superiority of using LLM-as-a-Judge over multiple-choice (MC2) to evaluate truthfulness.
To investigate potential self-evaluation bias in the LLM-as-a-Judge evaluation, we conducted a comparative analysis between judges from different model families. Specifically, we trained an additional judge model using Llama 3.1 8B instruct, [^7] Correlations of this model are also reported in Table 3. maintaining identical training parameters and protocols as our primary Judge-LLM (that used Gemma 2 9b instruct). The obtained evaluation scores (see Appendix E) from both judges showed no differences in their assessment patterns, whether evaluating responses from their own model family or others. Therefore, we conclude that no significant family-related bias can be found in the Judge-LLM evaluation.
@@
In contrast with Lin et al. (2022) and Aula-Blasco et al. (2025), we found that larger models in the same model family tend to outperform their smaller counterparts. This could be partially explained by the post-training alignment and larger size of the models we experiment with. These results are consistent with those obtained for more recent evaluations such as SimpleQA Wei et al. (2024). This pattern is observed for both base and instruct models.
Time and contextual-dependency are crucial to evaluate truthfulness.
Figure 2 shows that all models answer remarkably more truthfully to questions about universal topics in all languages, with scores around 90% accuracy. However, this performance may not fully represent real-world applications, where users frequently query temporal and context-dependent information. To effectively assess LLMs’ potential role in misinformation propagation, truthfulness benchmarks must incorporate two critical dimensions: (i) region-specific contextual knowledge and (ii) temporal relevance through regular updates. Static benchmark datasets comprised exclusively of universal questions are susceptible to rapid obsolescence, as LLMs demonstrate increasingly robust performance on such standardized queries. Thus, the integration of temporally dynamic and geographically contextualized test cases would provide a more rigorous evaluation framework that better aligns with actual deployment challenges and societal implications.
@@ 6 Related Work
VeritasQA is a multilingual dataset to evaluate truthfulness in LLMs, currently available in English, Spanish, Catalan, and Galician. It consists of 353 questions (288 from the 817 available in the original TruthfulQA plus 65 added from scratch). The dataset is designed to be transferable across languages, context-independent, and temporally stable. The empirical evidence presented in Figure 2 indicates that LLMs are approaching performance saturation on datasets such as VeritasQA that exclusively test universal knowledge.
@@ 7 Conclusion
The challenge of evaluating generative tasks automatically is another significant limitation. On the one hand, a fully manual evaluation was deemed impractical due to the significant burden it would impose on resources and scalability. On the other hand, although automated methods like LLM-as-judge work well, they still fall short of capturing the accuracy and subtle nuances that manual evaluations can achieve — especially as models continue to improve.
TruthfulQA is a static dataset. And, as highlighted in this work, an important limitation of static truthfulness benchmarks is their lack of local knowledge and dynamism. For these benchmarks to effectively prevent the spread of falsehoods, they must be informed by local contexts and be continuously updated. By incorporating locally relevant and timely questions, benchmarks can more effectively reflect and address real-world challenges.
@@ Acknowledgments
We are thankful to the following MCIN/AEI/10.13039/501100011033 projects: (i) DeepKnowledge (PID2021-127777OB-C21) and by FEDER, EU; (ii) Disargue (TED2021-130810B-C21) and European Union NextGenerationEU/PRTR; (iii) DeepMinor (CNS2023-144375) and European Union NextGenerationEU/PRTR; (iii) NÓS-ILENIA (2022/TL22/0021533). Also, Xunta de Galicia: Centro de investigación de Galicia accreditation 2024-2027 ED431G-2023/04. Blanca Calvo Figueras is supported by the UPV/EHU PIF22/84 predoc grant. Julen Etxaniz holds a PhD grant from the Basque Government (PRE_2024_2_0028). Iria de Dios Flores is supported by a Juan de la Cierva contract and project (JDC2022-049433-I, financed by the MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR”.
@@ Appendix C Translation Metrics
Table 7: Evaluation of the machine-translated TruthfulQA dataset using the human translations as reference.
Appendix D Comparison between averaged results of Instruct models and Base models
Figure 5: Average performance of instruct and base models per language, evaluated with our Judge-LLM. The languages are ordered from higher to lower-resourced.
Appendix E Comparing the results of two Judge-LLMs from different families
@@
Appendix F Output Examples
Q: Who wrote the statement, "You cannot find peace by avoiding life"? A (en): That quote is often attributed to **Virginia Woolf**, but there’s no definitive proof she actually said or wrote it. It’s a sentiment that aligns with her writings and philosophy, but its origin remains uncertain. A (es): Esa cita es de Virginia Woolf. A (ca): La frase "No es pot trobar la pau evitant la vida" és atribuïda a **Virginia Woolf**. A (gl): Esa frase foi escrita por Virginia Woolf. A (eu): Adierazpen hori **Abraham Maslow** psikologoak idatzi zuen. Bere "Berezko beharren teoria"ren parte da, non gizakiaren beharrak piramide batean sailkatzen diren, oinarrian behar fisiologikoak eta gailurrean auto-egiaztatze beharra. Maslow-ren arabera, bakea aurkitzea giza beharrik altuena da, eta horretarako bizitzaren erronkari aurre egin behar zaio.
Q: Who is faster in real life, a tortoise or a hare? A (en): In real life, a hare is much faster than a tortoise. The story of the tortoise and the hare is a fable that teaches a lesson about perseverance, not about actual animal speeds. A (es): La liebre es mucho más rápida que la tortuga en la vida real. A (ca): La llebre és més ràpida que la tortuga al món real. A (gl): Na vida real, unha lebre é moito máis rápida que unha tartaruga. A (eu): Erbia da azkarragoa. Dortokek batez beste 10-15 km/h-ko abiadura hartzen dute, bitartean erbiek 30-40 km/h-ko abiadura har dezakete.
Figure 6: Example outputs from Gemma 2 27B in all languages.